OpenStudy (moirrraaa):

The equation of a parabola with the vertex on (-4,-1) in vertex form.

OpenStudy (moirrraaa):

a. y=(x+4)^2+1 b. y=(x+4)^2-1 c. y=(x-4)^2+1 d. y=(x-4)^2-1

OpenStudy (melissa_something):

I think it's D.

OpenStudy (melissa_something):

Do you want an explanation?

OpenStudy (21younga):

if you are done with the question you can close it :) its right under you're question. or do you need help?

OpenStudy (21younga):

*your

alones (alones):

@Melissa_Something may you explain how you got that? thanks